OpenStudy (anonymous):

If x3 = -27, then what is x?

geerky42 (geerky42):

\(\large x^3 = -27 \Rightarrow \sqrt[3]{x^3} = x = \sqrt[3]{-27}\) \(\large\sqrt[3]{-27} = \space ?\)

geerky42 (geerky42):

HINT: 3³ = 27

OpenStudy (anonymous):

can be written as x^3+27=0 x^3+3^3=0 x^3+3^3=0 since \[\Large a^3+b^3=(a+b)(a^2-ab+b^2)\] (x+3)(x^2-3x+9)=0 x+3=0 x^2-3x+9=0 (use quaadratic equation you will get complex roots )